OpenStudy (anonymous):

Zach borrowed $4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved $5,440 to pay back to his grandfather. At what interest rate did Zach borrow the money? A) 9% B) 12% C) 25% D) 26%

OpenStudy (johnweldon1993):

So again here we have Interest earned...is current balance - original balance 5440 - 4000 = 1440....this is 'I' 't' is 3 years and P = 4000 so we want to solve for 'r' I = prt \[R = \frac{ I }{ pt }\] plug in your numbers \[R = \frac{ 1440 }{ 4000(3) }\] what does 'r' equal? :P